Palencia Paints Corporation has a target capital structure of 30% debt and 70% common equity, with no preferred stock. Its before-tax cost of debt is 11%, and its marginal tax rate is 25%. The current stock price is P0 = $33.00. The last dividend was D0 = $2.00, and it is expected to grow at a 6% constant rate. What is its cost of common equity and its WACC?
Answer #1

We can find the cost of common equity using "Dividend Growth method"

ke=D+9

D1 = D0 + g

=2 + 6% of 2

= 2.12

2.12 ::Ke = +0.06 33

=0.1242

or 12.42%

WACC = (Percentage of Equity * Cost of equity) + [Percentage of Debt * Before Tax Cost of Debt ( 1- Tax Rate)]

:: WACC = (0.70 * 12.42) + 0.30 * 11(1 - 25)

= (0.0870) + 0.02475

= 0.1117

or 11.17%